# Minimizing resource overhead in fusion-based quantum computation using hybrid spin-photon devices

### v1

@@ IV Discussion
Using low-loss fibers and interleaving, it is estimated that one RSG can effectively serve as several logical qubits in a surface code, leading to industrial-scale applications requiring between $10^{3}$ and $10^{5}$ RSGs [5]. Thus, with just 12 sources per RSG, an FBQC architecture using RUS-based RSGs could reach industrial-scale fault-tolerant computations using as few as 12000 sources. This is in stark contrast to a standard surface code with superconducting circuits that would require over a million physical qubits and also in contrast to the all-photonic FBQC architecture using HSPSs studied in this work that we estimate would require roughly 10 billion sources (6 million times more than a RUS-based architecture). Thus a hybrid spin-photonic approach can reduce the resource overhead of fault-tolerant quantum computation by multiple orders-of-magnitude compared to both all-matter and all-photonic platforms. Such a major reduction is expected to translate directly into a substantially reduce cost of fabrication and operation.
@@
For instance, it is possible to design photonic schemes using fewer stages of multiplexing, provided that switching networks are highly transmissive but where the majority of loss occurs due to the photonic memory needed to reconfigure the switching network. It may also be possible to determine a more optimal division of the construction into intermediate states or use of auxiliary resources [57].
@@
In addition to the optical depth, it is crucial to consider the efficiency of the source itself. For heralded sources, free-space implementations have shown efficiency values up to 94% [53] before lossy collection into a fiber. On-chip implementations necessary to mass fabricate sources can produce high purity photons, but reaching high source efficiency remains a challenge with values around 26% [54]. Deterministic sources have improved significantly over the past decade, with micropillar cavities consistently demonstrating efficiency values above $50\%$ before lossy collection into a fiber [28, 58]. On-chip sources have reached efficiencies up to 86% [59], but efficient coupling to low-loss waveguides or fibers has yet to be demonstrated. State-of-the-art values are obtained using a tunable open cavity design. This latter approach has demonstrated the highest fibered efficiencies, with $57\%$ in 2020 [52] and $71\%$ in 2023 [30]. However, the manufacturability and stability of open cavity designs remains challenging.
@@
In addition to photon collection efficiency, photons that are highly pure and indistinguishable are necessary to ensure high-fidelity fusion and RUS gates. Currently, HSPS hold the record for indistinguishability of up to 99.5% for photons from distinct sources, but come with a multi-photon error of 0.7% (integrated autocorrelation of $g^{(2)}(0)=0.35\%$) [54]. This also comes along with a low success probability (est. $p_{s}<1\%$) limited by spectral filters that cause $20\%$ loss [54] and hence prevent efficient photon-number resolved heralding that suppresses multi-photon emission produced at higher success probabilities. The state-of-the-art for indistinguishability between photons from distinct QDs in bulk material is 93.0% [60], but is expected to reach values exceeding $99\%$ when using a cavity to suppress phonon side-band emission [61]. Quantum dots also consistently provide sub-1% multi-photon errors without compromising other metrics [58], and they hold the record for multi-photon errors as low as 0.02% [62].
@@
On-chip optical components are also reaching the high levels of quality needed to implement linear-optical processing. Twenty-mode universal SiN chips are commercially available that reach $>97\%$ transformation fidelity [63]. Using machine learning characterization, universal interferometers can also reach fidelity exceeding 99.77% [64]. Smaller specialized integrated circuits show even higher fidelity, with a standard Mach-Zehnder interferometer reaching a fidelity up to 99.996% when characterized with coherent light [65, 54]. Component losses are also improving, with SiN integrated beam splitters and mode crossings reaching 0.012% and 0.037%, respectively [54]. The components that must still be improved are the integrated PNR detectors, the chip-to-fiber interface and high-bandwidth phase shifters. The last two currently incur a loss of 1.3% and 2.3%, respectively [54], which exceed the maximum loss per component estimated for all but the RUS-based architecture.
Solid-state QD sources with a matter qubit degree of freedom have recently demonstrated the generation of small entangled states [36, 37, 38, 39, 17]. These early demonstrations use devices with limited spin coherence times on the order of nanoseconds, which limits graph state sizes. However, it is predicted that 30-photon graph states with $>50\%$ fidelity can be achieved with only minor adjustments [17]. Moreover, state-of-the-art spin coherence times of up to 100 microseconds have been achieved with QDs using dynamical decoupling [66]. This is more than sufficient to construct a 24-photon resource state with $>99.95\%$ fidelity, even when limited by current feedforward times of 23 ns [67]. With this coherence time, resource state fidelity will be primarily limited by single-qubit gates, which can already reach 99.3% [66] and are expected to improve substantially in an industrial setting, in addition to the RUS gate fidelity, which depends primarily on the purity and indistinguishability of the emitted photons.

### v2

@@ IV Discussion
Using low-loss fibers and interleaving [63], it has been estimated in Ref. [5] that one RSG can effectively feed multiple logical qubits in the fusion network, enabling industrial-scale applications with between $10^{3}$ and $10^{5}$ RSGs. This benefit applies broadly to FBQC architectures, regardless of the specific RSG implementation. Thus, with a minimum of 12 sources per RSG, an FBQC architecture using RUS-based RSGs could achieve industrial-scale fault-tolerant computation with as few as 12000 sources. This is in stark contrast to a surface code architecture implemented with superconducting circuits, which would require over a million physical qubits, and to the all-photonic FBQC architecture using HSPSs studied in this work, which we estimate would require upwards of one billion sources—around five orders of magnitude more than a RUS-based architecture. A hybrid spin-photonic approach can therefore reduce the resource overhead of fault-tolerant quantum computation by multiple orders of magnitude compared to both all-matter and all-photonic platforms. Such a major reduction is expected to translate directly into a substantially lower fabrication and operation costs.
@@
For instance, it is possible to design photonic schemes using fewer stages of multiplexing, provided that switching networks are highly transmissive but where the majority of loss occurs due to the photonic memory needed to reconfigure the switching network. It may also be possible to determine a more optimal division of the construction into intermediate states [55], better use of auxiliary resources [64], or the recycling of photonic states after failed operations [65] at the cost of increased loss and error.
@@
Our analysis provides a foundation for extending comparisons to other graph codes and construction schemes, enabling a broader mapping of RSG architectures in terms of resource efficiency and optical depth for different source types. While we have focused on critical metrics—component count, entanglement success, and loss tolerance—that have a profound impact, refined comparisons must also incorporate other errors that arise during state construction and extend the analysis to full error correction simulations in order to determine fault-tolerance thresholds. Although accurate threshold simulations are beyond the scope of this work, in Appendix J, we estimate the number of effective photon-photon and spin-photon interactions required by each scheme. This analysis reveals that gains in resource efficiency are generally accompanied by an increased reliance on spin-photon interactions. Notably, while the balance between photon-photon and spin-photon operations varies across schemes, the total number of effective two-qubit interactions remains broadly comparable. This suggests that, at a first approximation, all four schemes may exhibit similar overall sensitivity to noise. However, capturing the true performance impact will require detailed modeling of physically motivated error channels and full-scale simulations of the noisy construction circuits.
@@
In addition to the optical depth, it is crucial to consider the efficiency of the source itself. For heralded sources, free-space implementations have shown efficiency values up to 94% [59] before lossy collection into a fiber. On-chip implementations necessary to mass fabricate sources can produce high purity photons, but reaching high source efficiency remains a challenge with values around 26% [60]. Deterministic sources have improved significantly over the past decade, with micropillar cavities consistently demonstrating efficiency values above $50\%$ before lossy collection into a fiber [28, 66]. On-chip sources have reached efficiencies up to 84% [67], but efficient coupling to low-loss waveguides or fibers has yet to be demonstrated. State-of-the-art values are obtained using a tunable open cavity design. This latter approach has demonstrated the highest fibered efficiencies, with $57\%$ in 2020 [58] and $71\%$ in 2023 [30]. However, the manufacturability and stability of open cavity designs remains challenging.
@@
In addition to photon collection efficiency, photons that are highly pure and indistinguishable are necessary to ensure high-fidelity fusion and RUS gates. Currently, HSPS hold the record for indistinguishability of up to 99.5% for photons from distinct sources, but come with a multi-photon error of 0.7% (integrated autocorrelation of $g^{(2)}(0)=0.35\%$) [60]. This also comes along with a low success probability (est. $p_{s}<1\%$) limited by spectral filters that cause $20\%$ loss [60] and hence prevent efficient photon-number resolved heralding that suppresses multi-photon emission produced at higher success probabilities. The state-of-the-art indistinguishability between photons emitted from distinct quantum dots in bulk material is 93.0% [68], limited by a combination of phonon-induced dephasing and charge noise. Further improvements are expected by combining the high-quality fabrication of Ref. [68] with cavity-enhanced emission [69], which suppresses the phonon sideband and shortens the emission lifetime—reducing both dephasing of the zero-phonon line and charge noise to enable indistinguishability above 99%. Quantum dots also consistently provide sub-1% multi-photon errors without compromising other metrics [66], and they hold the record for multi-photon errors as low as 0.02% [70].
@@
On-chip optical components are also reaching the high levels of quality needed to implement linear-optical processing. Twenty-mode universal SiN chips are commercially available that reach $97\%$ transformation fidelity [71]. Using machine learning characterization, universal interferometers can also reach fidelity exceeding 99.77% [72]. Smaller specialized integrated circuits show even higher fidelity, with a standard Mach-Zehnder interferometer reaching a fidelity up to 99.996% when characterized with coherent light [73, 60]. Component losses are also improving, with SiN integrated beam splitters and mode crossings reaching 0.012% and 0.037%, respectively [60]. The components that must still be improved are the integrated PNR detectors, the chip-to-fiber interface and high-bandwidth phase shifters. The last two currently incur a loss of 1.3% and 2.3%, respectively [60], which exceed the maximum loss per component estimated for all but the RUS-based architecture.
[…]
With extended coherence times and improved spin-photon entanglement quality, resource state fidelity will be primarily limited by single-qubit gates, which can already reach 99.3% [74] and are expected to improve substantially in an industrial setting, in addition to the RUS gate fidelity, which depends primarily on the purity and indistinguishability of the emitted photons.